OpenStudy (anonymous):

How would I solve this? (a/b+1/a) (a-(a3)/(a2-b))

OpenStudy (anonymous):

one of those is the answer

OpenStudy (anonymous):

So, solve what's in the parentheses first, then multiply them.

OpenStudy (anonymous):

What exactly are you having trouble with?

OpenStudy (anonymous):

So the first parentheses is (a2+b)/(ab) and the second one is (-ab)/(a2-b) ?

OpenStudy (anonymous):

One second

OpenStudy (anonymous):



OpenStudy (anonymous):

Now, subtract the a^3 and you'll have ab. So, \[equation1(\frac{ a ^{2}+b }{ ab}) (\frac{ ab }{ a ^{2}+b })equation2\] Now, notice that equation 1 has as its denominator ab, and equation 2 has as its numerator ab.... cross them out and replace them with 1 :D \[equation1(\frac{ a ^{2}+b }{ 1})(\frac{ 1}{a ^{2}+b})\] Which, equals \[\frac{ a ^{2}+b }{ a ^{2}+b}\] =1!

OpenStudy (anonymous):

Thank you!